Evaluate the integral. tan 5x sec25x dx al ſ tan 5x sec ?5x dx = 0 TEL
Answer #1

L = Stan(5x) * sec (5x) dx Let u = tan(5x) On differentiating, du = sec (5x) * 5 dx {Derivative of tan(ax) = sec (ax) * a} du